the width of a rectangle is 2m less than the length.the area is 80cm square,find the lengths of the base and width
Answer by Felixozabs(1)

Let length =y
let width =x
recall Y x X=80
x=y-2 i.e 2cm less than y
y x (y-2)=80
y^2 -2y=80
y^2 -2y-80=0
(y-10)(y+8)=0
therefore y=10
recall that x=(y-2)=80
if y=10 then x=10-2=8
length=10
width=8